Sailor:

I am absolutely stuck on this: It's Geometry, you have to write the equation in standard form (x-h)^2+(y-k)^2=r^2

Sailor:

So you have to use the center (-3,-2) and the point on the circle (5,7)

surjithayer:

radius=point on the circle-center of the circle. \[r=\sqrt{(5+3)^2+(7+2)^2}=\sqrt{64+81}=\sqrt{145}\] (h,k)=(-3,-2) h=-3 k=-2 eq. of circle is \[(x+3)^2+(y+2)^2=(\sqrt{145})^2\] \[x^2+6x+9+y^2+4y+4=145\] simplify and write in the form\[x^2+y^2+2gx+2fy+c=0\]